OpenStudy (anonymous):

y = 3x -1 is graphed. you have a parallel line that gose thru point 2,0. what is the equation for the other line.

OpenStudy (freethinker):

parallel lines are lines of the same slope

OpenStudy (freethinker):

use the point slope form using the same slope as given with the points it goes through

OpenStudy (esshotwired):

If you know point-slope, then you can plug everything in and solve. So with this one it would be: y-0=3(x-2)

OpenStudy (freethinker):

oops

OpenStudy (freethinker):

so knowing that your slope is 3 from the first y-intercept form equation of y=3x-1 you can use it in the point-slope form to look for the equation of the other line\[y-y_1=m(x-x_1)\]

OpenStudy (esshotwired):

Then you would add 0 to both sides(nothing will change). Then distribute. After you distribute, you will get the answer. Try it and I will tell you if you are correct.